# Optimal Task Order for Continual Learning of Multiple Tasks

### v1

@@ 1 Introduction
The ability to learn multiple tasks continuously is a hallmark of general intelligence. However, deep neural networks and its applications, including large language models, struggle with continual learning and often suffer from catastrophic forgetting of previously acquired knowledge [1, 2, 3, 4]. Although extensive work has been done to identify when forgetting is most prevalent [5, 6] and how to mitigate it [7, 8, 9, 10, 11, 12], it remains unclear how to prevent forgetting while simultaneously promoting knowledge transfer across tasks [13, 14, 15, 16].
One important yet relatively underexplored aspect of continual learning is task-order dependence. Previous work has revealed that the order in which tasks are presented can significantly influence continual learning performance and also explored various approaches to optimize task order [17, 18, 19, 20, 21, 22]. However, we still lack clear understanding on how ordering of tasks influences the learning performance and how to order tasks to achieve optimal performance. Figure 1 illustrates this problem using a continual binary image classification example, where a neural network is trained on three tasks: A (Cat vs. Ship), B (Frog vs. Truck), and C (Horse vs. Deer). Learning one task can influence performance on the others in a complex manner (Figs. 1a and 1b). Consequently, in this example, the CBA task order achieves a higher average performance after training than the ABC order (Fig. 1c). The goal of this work is to understand this task-order dependence in continual learning.
@@
Task-order optimization requires data samples from all tasks beforehand, making it infeasible in a strictly online learning setting. Nevertheless, it remains highly relevant for many learning problems. For instance, if the current task creates unfavorable conditions, systems can postpone its learning to a more suitable time. In addition, understanding how the order of tasks impacts learning can serve as a tool for predicting the difficulty of online learning given a data stream. Moreover, even when all tasks are known to the learner a priori, training them jointly in an offline manner can be impractical or costly, making a continual learning with the optimal task-order a desirable alternative. For instance, large language models are typically trained in an online fashion because the size of the training corpus is so vast that multiple epochs of training over the entire dataset is infeasible [23, 24]. In such cases, how the corpus is organized for training can significantly impact learning efficiency, supporting the importance of optimizing task/corpus order. Furthermore, in robotics applications [25], transitioning from one task to another demands time and effort, making simultaneous multi-task learning inefficient. Similar constraints arise in designing of machine-learning-based teaching curricula for schools or professional training where learners need to study multiple subjects sequentially [26, 22].
@@
We then investigate how the similarity between tasks, when placed in various positions within the task order, affects the overall error. Through linear perturbation analysis, we find that the task-order effect decomposes into two factors. The first is absolute order dependence: similarity between two tasks influences the error differently depending on whether these tasks appear near the beginning or near the end of the sequence. We demonstrate that when tasks are on average positively correlated, the least representative tasks should be learned first, while the most typical task should be learned last (periphery-to-core rule). The second factor is relative order dependence: the effect of task similarity on the error differs depending on whether two tasks are adjacent in the sequence or far apart. We show that a task order maximizing the path length in the task dissimilarity graph outperforms one that minimizes this path length (max-path rule), consistent with previous empirical observations [20].
We illustrate these two rules by applying them to tasks with simple similarity structures forming chain, ring, and tree graphs, revealing the presence of non-trivial task orders that robustly achieve the optimal learning performance, given a graph structure. Moreover, we apply these rules to continual image classification tasks using the Fashion-MNIST, CIFAR-10, and CIFAR-100 datasets. We estimate task similarity by measuring the zero-shot transfer performance between tasks, and then implement the task-ordering rules based on these estimates. Our results show that both the periphery-to-core rule and the max-path rule hold robustly in both multilayer perceptrons and convolutional neural networks. This work thus provides a simple and generalizable theory to task-order optimization in task-incremental continual learning.
@@ 2 Related work
The effects of curriculum learning have been extensively studied in the reinforcement learning (RL) literature [27, 28, 29]. However, these studies primarily focus on learning a single challenging task by sequentially training on simpler tasks, leaving open the question of how to design a curriculum for learning multiple tasks of similar difficulty. A limited number of works have explored task-order optimization for continual/lifelong learning across multiple tasks by contrast.
Lad et al. [17] demonstrated that ordering tasks based on pairwise order preferences can lead to better classification performance compared to random task ordering. More recently, Bell and Lawrence [20] investigated task-order optimization by examining Hamiltonian paths on a task dissimilarity graph (see Sec. 4 for details). They hypothesized that the shortest Hamiltonian path would be optimal but instead found that the longest Hamiltonian path significantly outperformed both random task ordering and the shortest path in continual image classification tasks. Our work provides analytical insights into when and why this is the case. Lin et al. [21] analyzed generalization error and task-order optimization in continual learning for linear regression. Our work advances this theoretical framework in several important ways. First, we introduce a latent structure model for considering the effect of input similarity and reveal how tasks’ relative positions—not just their absolute positions as in Lin et al. [21]’s Equation 10—influence the model’s final performance. We validate this theoretical finding through experiments on both synthetic data and image classification tasks. Furthermore, we extend beyond the synthetic task settings of Lin et al. [21] by demonstrating these effects in a general continual learning framework using data-driven similarity estimation. Task-order effects on continual learning have also been analyzed in [18, 30, 22].
The linear teacher-student model used in this work is a widely adopted framework for analyzing the average properties of neural networks by explicitly modeling the data generation process through a teacher model [31, 32, 33]. Due to their analytical tractability, these models have offered deep insights into various aspects of statistical learning problems, including generalization [34, 35], learning dynamics [36, 37, 38], and representation learning [39, 40]. Many studies have also applied this framework to explore various aspects of continual learning [41, 6, 42, 43, 44, 21, 45, 46, 47].
@@ 3.1 Model setting
Here, $\bm{s}\in\mathbb{R}^{N_{s}}$ is the latent factor that underlies both $\bm{x}$ and $\bm{y}^{*}$, $I$ is the identity matrix, and $A_{\mu}\in\mathbb{R}^{N_{x}\times N_{s}}$ and $B_{\mu}\in\mathbb{R}^{N_{y}\times N_{s}}$ are the mixing matrices that generate the input $\bm{x}$ and the target $\bm{y}^{*}$ from the latent $\bm{s}$ (Fig. 2a). Below we focus on $N_{x}\gg N_{s}$ regime. The introduction of this low-dimensional latent factor $\bm{s}$ is motivated by the presence of low-dimensional latent structures in many real-world datasets [48, 49].
@@
where $C^{in}$ is a $P\times P$ matrices that specify input correlation between tasks. Similarly, we sample the ($i,j$)-th elements of $B_{1},..,B_{P}$, $\bm{b}_{ij}\equiv[B_{1,ij},B_{2,ij},....,B_{P,ij}]^{T}$, by $\bm{b}_{ij}\sim\mathcal{N}(\bm{0},\tfrac{1}{N_{s}}C^{out})$. Note that here correlation is introduced across tasks in an element-wise manner while keeping elements of each mixing matrix independent (i.e. $\left\langle A^{\mu}_{ij}A^{\nu}_{kl}\right\rangle_{A}=\delta_{ik}\delta_{jl}\tfrac{C^{in}_{\mu\nu}}{N_{s}}$). Here we generate the model from the task similarity matrices $\{C^{in},C^{out}\}$ because previous work suggests the crucial impact of task similarity on continual learning [5, 6]. In section 5, we consider the estimation of task similarity from datasets to ensure the applicability of our framework.
@@
where $\left\langle\cdot\right\rangle_{\bm{s}}$ represents expectation over latent $\bm{s}\sim\mathcal{N}(\bm{0},I)$. In a task-incremental continual learning task [50], we are mainly concerned with minimizing the total error on all tasks after learning all tasks. Denoting the network parameter after learning of the last, $P$-th, task by $W_{P}$, the final error is defined by
@@ 4.2 Impact of task typicality
If $01>m>$ 0, $\tfrac{3-m}{(2-m)(1-m)}>0$ and $(1-m)^{\mu}\geq(1-m)^{P}>\alpha^{+}$ for $\mu=1,2,...,P$. Therefore, $G^{+}_{\mu\nu}$ is a monotonically decreasing function of both $\mu$ and $\nu$ under $01>m>$ 0 (Fig. 3d). This means that, to minimize the error contributed from $G^{+}_{\mu\nu}$, $\delta\epsilon^{+}_{f}\equiv\sum_{\mu,\nu}G^{+}_{\mu\nu}\delta M_{\mu\nu}$, the tasks should be ordered in a way that the residual similarity $\delta M_{\mu\nu}$ takes a small (preferably negative) value for early task pairs and a large value for later task pairs. In other words, earlier pairs should be relatively dissimilar to each other, while later pairs should be mroe similar.
One heuristic way to achieve this task order is to put the most atypical task at the beginning and the most typical one at the end. Denoting the relative typicality of the task by $\delta t_{\mu}=\sum_{\nu\neq\mu}\delta M_{\nu\mu}$, if we arrange tasks as $\delta t_{1}\leq\delta t_{2}\leq...\leq\delta t_{P}$, on average, earlier pairs are dissimilar to each other while the latter ones are similar. Below, we denote this ordering as a periphery-to-core rule, as less representative periphery tasks are learned first and more central core tasks are learned later under this principle (Fig. 4b). Under a randomly generated input correlation matrix $C^{in}$, periphery-to-core task order robustly outperformed both random and core-to-periphery order, when the average correlation is large positive value (blue vs black and red line in Fig. 3e). This was not the case when the average correlation is a small positive value potentially due to contribution from $G^{-}$ factor. Note that, a similar rule was derived by Lin et al. [21] based on their analysis of linear regression model, where they proved that when there is one outlier task, the outlier task should be learned in the first half of the task sequence.
@@ 4.3 Impact of Hamiltonian path length
Given a set of tasks, let us define a task dissimilarity graph by setting each task as a node and dissimilarity between two tasks as the weight of the edge between corresponding nodes (Fig. 4c). Then, a task order that learns each task only once forms a Hamiltonian path on the graph, a path that visits all nodes once but only once. We can then construct a heuristic solution for minimizing $\delta\epsilon^{-}_{f}$ by selecting a task order that yields the longest Hamiltonian path. When tasks have the same similarity with each other in terms of $C^{out}$, their similarity depends solely on $C^{in}$, allowing us to define dissimilarity as $d_{\mu\nu}\equiv 1-C^{in}_{\mu\nu}$. Thus, the total length of the Hamiltonian path induced by a given task order follows $D_{H}=\sum_{\mu=1}^{P-1}C^{in}_{\mu,\mu+1}$. Consequently, $\delta\epsilon^{-}_{f}$ is rewritten as
$\epsilon^{-}_{f}$ $=-\alpha_{-}\left((1-m)^{P-1}D_{H}+\sum_{d=2}^{P-1}(1-m)^{P-d}\sum_{\mu=1}^{P-d}d_{\mu\nu}\right)$
[…]
Because $-\alpha_{-}$ is non-negative, small task dissimilarity $d_{\mu\nu}$ (i.e., large task correlation $C^{in}_{\mu\nu}$) generally helps minimizing the error. Moreover, we have $0\leq(1-m)^{P-1}<(1-m)^{P-d}$ for $d=2,3,...$, indicating $D_{H}$ term has the smallest impact on the error. Therefore, by choosing the largest $d_{\mu\nu}$ for $D_{H}$, we can make $\delta\epsilon^{-}_{f}$ small on average. We observed this trend robustly even when we sampled $\{C^{in},C^{out}\}$ randomly (Fig. 3g). Our work thus provides theoretical insights on why the maximum Hamiltonian path provides a preferable task order, strengthening previous empirical finding [20]. We call this rule as max-path rule below.
@@ 5.1 Empirical estimation of task similarity
To apply these principles, it is necessary to first measure the similarity between tasks. Here, we estimate the similarity between tasks $A$ and $B$ by measuring the zero-shot transfer performance between them (Fig. 8 in the Appendix). Specifically, we train a network for task $A$, obtaining the learned weights $W_{A}$. We then measure the error of this trained network on task $B$, denoted as $\epsilon_{B}[W_{A}]$. Since the transfer performance from task $A$ to $B$ generally differs from that of $B$ to $A$  we take the mean of both directions and define the similarity $\rho_{AB}$:

### v2

@@ 1 Introduction
The ability to learn multiple tasks continuously is a hallmark of general intelligence. However, deep neural networks and its applications, including large language models, struggle with continual learning and often suffer from catastrophic forgetting of previously acquired knowledge (McCloskey & Cohen, 1989; French, 1999; Hadsell et al., 2020; Luo et al., 2023). Although extensive work has been done to identify when forgetting is most prevalent (Ramasesh et al., ; Lee et al., 2021) and how to mitigate it (French, 1991; Robins, 1995; Kirkpatrick et al., 2017; Shin et al., 2017; Serra et al., 2018; Rolnick et al., 2019), it remains unclear how to prevent forgetting while simultaneously promoting knowledge transfer across tasks (Ke et al., 2020; Lin et al., 2022; Ke et al., ; Kontogianni et al., 2024).
One important yet relatively underexplored aspect of continual learning is task-order dependence. Previous work has revealed that the order in which tasks are presented can significantly influence continual learning performance and also explored various approaches to optimize task order (Lad et al., 2009; Pentina et al., 2015; Guo et al., 2018; Bell & Lawrence, 2022; Lin et al., 2023; Singh et al., 2023). However, we still lack clear understanding on how ordering of tasks influences the learning performance and how to order tasks to achieve optimal performance. Figure 1 illustrates this problem using a continual binary image classification example, where a neural network is trained on three tasks: A (Cat vs. Ship), B (Frog vs. Truck), and C (Horse vs. Deer). Learning one task can influence performance on the others in a complex manner (Figs. 1a and 1b). Consequently, in this example, the CBA task order achieves a higher average performance after training than the ABC order (Fig. 1c). The goal of this work is to understand this task-order dependence in continual learning.
@@
Task-order optimization requires some amount of knowledge on all tasks beforehand, making it infeasible in a strictly online learning setting. Nevertheless, it remains highly relevant for many learning problems. One scenario is when data acquisition and training need to be conducted in parallel, which may occur in the training of self-driving algorithms (Verwimp et al., 2023) or medical image analysis (Kumari et al., 2023). In this setting, it is beneficial to first collect a small pilot dataset across all underlying tasks and then determine the optimal order of data acquisition and training to maximize knowledge transfer while minimizing forgetting across tasks. To demonstrate the potential applicability of our theory to this problem, we provide numerical evidence that, in continual visual recognition benchmarks, an optimal task order estimated from just 1% of the data significantly outperforms a random task order.
Moreover, in robotics applications (Lesort et al., 2020; Ibarz et al., 2021), switching between tasks often involves physically rearranging objects around the robot, which is both time-consuming and labor-intensive. As a result, switching tasks on a trial-by-trial basis is often infeasible, necessitating block-wise training. In this scenario, optimizing task order could help maximize average performance across all tasks. Similar constraints arise in designing of machine-learning-based teaching curricula for schools or professional training where learners need to study multiple subjects sequentially (Rafferty et al., 2016; Singh et al., 2023). Furthermore, even in a more traditional continual learning task, if the current task creates unfavorable conditions, systems can postpone its learning to a more suitable time. Understanding how the order of tasks impacts learning can also serve as a tool for predicting the difficulty of online learning given a data stream. Lastly, large language models are typically trained in an online fashion because the size of the training corpus is so vast that multiple epochs of training over the entire dataset is infeasible (Hoffmann et al., 2022; Chowdhery et al., 2023). In such cases, how the corpus is organized for training can significantly impact learning efficiency, supporting the importance of optimizing task/corpus order.
@@
We then investigate how the similarity between tasks, when placed in various positions within the task order, affects the overall error. Through linear perturbation analysis, we find that the task-order effect decomposes into two factors. The first is absolute order dependence: similarity between two tasks influences the error differently depending on whether these tasks appear near the beginning or near the end of the sequence. We demonstrate that when tasks are on average positively correlated, the least representative tasks should be learned first, while the most typical task should be learned last (periphery-to-core rule). The second factor is relative order dependence: the effect of task similarity on the error differs depending on whether two tasks are adjacent in the sequence or far apart. We show that a task order maximizing the path length in the task dissimilarity graph outperforms one that minimizes this path length (max-path rule), consistent with previous empirical observations (Bell & Lawrence, 2022).
We illustrate these two rules by applying them to tasks with simple similarity structures forming chain, ring, and tree graphs, revealing the presence of non-trivial task orders that robustly achieve the optimal learning performance, given a graph structure. Moreover, we apply these rules to continual image classification tasks using the Fashion-MNIST, CIFAR-10, and CIFAR-100 datasets. We estimate task similarity by measuring the zero-shot transfer performance between tasks, and then implement the task-ordering rules based on these estimates. Our results show that both the periphery-to-core rule and the max-path rule hold robustly in both multilayer perceptrons and convolutional neural networks. Moreover, using $\sim 1\%$ of the data for the task similarity and order estimation was sufficient to achieve a significant improvement over random ordering. This work thus provides a simple and generalizable theory to task-order optimization in task-incremental continual learning.
@@ 2 Related Work
The effects of curriculum learning have been extensively studied in the reinforcement learning (RL) literature (Elman, 1993; Krueger & Dayan, 2009; Narvekar et al., 2020). However, these studies primarily focus on learning a single challenging task by sequentially training on simpler tasks, leaving open the question of how to design a curriculum for learning multiple tasks of similar difficulty. A limited number of works have explored task-order optimization for continual/lifelong learning across multiple tasks by contrast.
Lad et al. (2009) demonstrated that ordering tasks based on pairwise order preferences can lead to better classification performance compared to random task ordering. More recently, Bell & Lawrence (2022) investigated task-order optimization by examining Hamiltonian paths on a task dissimilarity graph (see Sec. 4 for details). They hypothesized that the shortest Hamiltonian path would be optimal but instead found that the longest Hamiltonian path significantly outperformed both random task ordering and the shortest path in continual image classification tasks. Our work provides analytical insights into when and why this is the case. Lin et al. (2023) analyzed generalization error and task-order optimization in continual learning for linear regression. Our work advances this theoretical framework in several important ways. First, we introduce a latent structure model for considering the effect of input similarity and reveal how tasks’ relative positions—not just their absolute positions as in Lin et al. (2023)’s Equation 10—influence the model’s final performance. We validate this theoretical finding through experiments on both synthetic data and image classification tasks. Furthermore, we extend beyond the synthetic task settings of Lin et al. (2023) by demonstrating these effects in a general continual learning framework using data-driven similarity estimation. Task-order effects on continual learning have also been analyzed in (Pentina et al., 2015; Evron et al., 2023; Singh et al., 2023).
@@
The linear teacher-student model used in this work is a widely adopted framework for analyzing the average properties of neural networks by explicitly modeling the data generation process through a teacher model (Gardner & Derrida, 1989; Zdeborová & Krzakala, 2016; Bahri et al., 2020). Due to their analytical tractability, these models have offered deep insights into various aspects of statistical learning problems, including generalization (Seung et al., 1992; Advani et al., 2020), learning dynamics (Saad & Solla, 1995; Werfel et al., 2003; Saxe et al., 2014), and representation learning (Saxe et al., 2019; Tian et al., 2021). Many studies have also applied this framework to explore various aspects of continual learning (Asanuma et al., 2021; Lee et al., 2021; Evron et al., 2022; Goldfarb & Hand, 2023; Li et al., 2023; Lin et al., 2023; Goldfarb et al., ; Hiratani, ; Mori et al., ).
@@ 3.1 Model Setting
Here, $\bm{s}\in\mathbb{R}^{N_{s}}$ is the latent factor that underlies both $\bm{x}$ and $\bm{y}^{*}$, $I$ is the identity matrix, and $A_{\mu}\in\mathbb{R}^{N_{x}\times N_{s}}$ and $B_{\mu}\in\mathbb{R}^{N_{y}\times N_{s}}$ are the mixing matrices that generate the input $\bm{x}$ and the target $\bm{y}^{*}$ from the latent $\bm{s}$ (Fig. 2a). Below we focus on $N_{x}\gg N_{s}$ regime. The introduction of this low-dimensional latent factor $\bm{s}$ is motivated by the presence of low-dimensional latent structures in many real-world datasets (Yu et al., 2017; Cohen et al., 2020).
@@
where $C^{in}$ is a $P\times P$ matrices that specify input correlation between tasks. Similarly, we sample the ($i,j$)-th elements of $B_{1},..,B_{P}$, $\bm{b}_{ij}\equiv[B_{1,ij},B_{2,ij},....,B_{P,ij}]^{T}$, by $\bm{b}_{ij}\sim\mathcal{N}(\bm{0},\tfrac{1}{N_{s}}C^{out})$. Note that here correlation is introduced across tasks in an element-wise manner while keeping elements of each mixing matrix independent (i.e. $\left\langle A^{\mu}_{ij}A^{\nu}_{kl}\right\rangle_{A}=\delta_{ik}\delta_{jl}\tfrac{C^{in}_{\mu\nu}}{N_{s}}$), where $\delta_{ik}$ represents the Kronecker delta. Here we generate the model from the task similarity matrices $\{C^{in},C^{out}\}$ because previous work suggests the crucial impact of task similarity on continual learning (Ramasesh et al., ; Lee et al., 2021). In section 5, we consider the estimation of task similarity from datasets to ensure the applicability of our framework.
@@
where $\left\langle\cdot\right\rangle_{\bm{s}}$ represents expectation over latent $\bm{s}\sim\mathcal{N}(\bm{0},I)$. In a task-incremental continual learning task (Van de Ven & Tolias, 2019), we are mainly concerned with minimizing the total error on all tasks after learning all tasks. Denoting the network parameter after learning of the last, $P$-th, task by $W_{P}$, the final error is defined by
@@ 4.2 Impact of Task Typicality
If $01>m>$ 0, $\tfrac{3-m}{(2-m)(1-m)}>0$ and $(1-m)^{\mu}\geq(1-m)^{P}>\alpha^{+}$ for $\mu=1,2,...,P$. Therefore, $G^{+}_{\mu\nu}$ is a monotonically decreasing function of both $\mu$ and $\nu$ under $01>m>$ 0 (Fig. 3d). This means that, to minimize the error contributed from $G^{+}_{\mu\nu}$, $\delta\epsilon^{+}_{f}\equiv\sum_{\mu,\nu}G^{+}_{\mu\nu}\delta M_{\mu\nu}$, the tasks should be ordered in a way that the residual similarity $\delta M_{\mu\nu}$ takes a small (preferably negative) value for early task pairs and a large value for later task pairs. In other words, earlier pairs should be relatively dissimilar to each other, while later pairs should be more similar.
One heuristic way to achieve this task order is to put the most atypical task at the beginning and the most typical one at the end. Denoting the relative typicality of the task by $\delta t_{\mu}=\sum_{\nu\neq\mu}\delta M_{\nu\mu}$, if we arrange tasks as $\delta t_{1}\leq\delta t_{2}\leq...\leq\delta t_{P}$, on average, earlier pairs are dissimilar to each other while the latter ones are similar. Below, we denote this ordering as a periphery-to-core rule, as less representative periphery tasks are learned first and more central core tasks are learned later under this principle (Fig. 4b). Under a randomly generated input correlation matrix $C^{in}$, periphery-to-core task order robustly outperformed both random and core-to-periphery order, when the average correlation is large positive value (red vs black and blue line in Fig. 3e). This was not the case when the average correlation is a small positive value potentially due to contribution from $G^{-}$ factor. Note that, a similar rule was derived by Lin et al. (2023) based on their analysis of linear regression model, where they proved that when there is one outlier task, the outlier task should be learned in the first half of the task sequence.
@@ 4.3 Impact of Hamiltonian Path Length
Given a set of tasks, let us define a task dissimilarity graph by setting each task as a node and dissimilarity between two tasks as the weight of the edge between corresponding nodes (Fig. 4c). Then, a task order that learns each task only once forms a Hamiltonian path on the graph, a path that visits all nodes once but only once. We can then construct a heuristic solution for minimizing $\delta\epsilon^{-}_{f}$ by selecting a task order that yields the longest Hamiltonian path. When tasks have the same similarity with each other in terms of $C^{out}$, their similarity depends solely on $C^{in}$, allowing us to define dissimilarity as $d_{\mu\nu}\equiv 1-C^{in}_{\mu\nu}$. Thus, the total length of the Hamiltonian path induced by a given task order follows $D_{H}=\sum_{\mu=1}^{P-1}d_{\mu,\mu+1}$. Consequently, $\delta\epsilon^{-}_{f}$ is rewritten as
$\delta\epsilon^{-}_{f}=-\alpha^{-}$ $\Big{(}(1-m)^{P-1}D_{H}$
$+\sum_{d=2}^{P-1}(1-m)^{P-d}\sum_{\mu=1}^{P-d}d_{\mu,\mu+d}\Big{)}+\text{const.}$. (15)
Because $-\alpha^{-}$ is non-negative, small task dissimilarity $d_{\mu\nu}$ (i.e., large task correlation $C^{in}_{\mu\nu}$) generally helps minimizing the error. Moreover, we have $0\leq(1-m)^{P-1}<(1-m)^{P-d}$ for $d=2,3,...$, indicating $D_{H}$ term has the smallest impact on the error. Therefore, by choosing the largest $d_{\mu\nu}$ for $D_{H}$, we can make $\delta\epsilon^{-}_{f}$ small on average. We observed this trend robustly even when we sampled $\{C^{in},C^{out}\}$ randomly (Fig. 3g). Our work thus provides theoretical insights on why the maximum Hamiltonian path provides a preferable task order, strengthening previous empirical finding (Bell & Lawrence, 2022). We call this rule as max-path rule below.
@@ 5.1 Empirical Estimation of Task Similarity
To apply these principles, it is necessary to first measure the similarity between tasks. Here, we estimate the similarity between tasks $A$ and $B$ by measuring the zero-shot transfer performance between them (Fig. 8). Specifically, we train a network for task $A$, obtaining the learned weights $W_{A}$. We then measure the error of this trained network on task $B$, denoted as $\epsilon_{B}[W_{A}]$. Since the transfer performance from task $A$ to $B$ generally differs from that of $B$ to $A$  we take the mean of both directions and define the similarity $\rho_{AB}$: